OpenStudy (anonymous):

Write 7.9625 as a mixed number in simplest form.

OpenStudy (anonymous):

you want to get rid of the decimal and turn it into a fraction: 7.9625 = 79625 / 10000 now reduce the fraction 79625 / 10000 = 3185 / 400 = 637 / 80

OpenStudy (anonymous):

you could then turn it into a mixed fraction: 637/80 = 7 and 77/80